# Stolz positive scalar curvature structure groups, proper actions and equivariant 2-types

### v1

@@ 4 Invariance of R-groups under 2-equivalence
Let $f\colon X\rightarrow Y$ be a continuous, 2-connected $G$-map between $G$-CW-complexes (in particular, it induces a bijection between components of fixed point sets and an isomorphism of the fundamental groups of all components of fixed point sets of $X$ and the corresponding ones of $Y$). Then the functorially induced map
@@
Let us consider a class $[W,\varphi\colon W\rightarrow Y,g]\in R_{n}^{spin}(Y)^{G}$. We want to find a bordant cycle whose reference map factors through $f$.
Consider W as a bordism between its boundary $\partial W$ and the empty set and choose a $G$-invariant Morse function $\alpha\colon W\to\mathbb{R}$ on it with critical points rearranged as described in [Milnor, Theorem 4.8], namely for any critical points $p_{i}$ and $p_{j}$ such that $f(p_{i})<f(p_{j})$, we have that $Ind(p_{i})<Ind(p_{j})$. Notice that we are going to use the enhanced version of this result to the equivariant setting, see for instance [Mayer] or [Wasserman].
Then there exists a suitable $t\in\mathbb{R}$ such that the subset $W_{1}:=\alpha^{-1}([0,t])\subset W$ consists only of $G$-handles of dimensions 0,1 and 2. We immediately obtain a decomposition of $W$ as $W_{1}\cup W_{2}$ such that $W_{1}$ is a bordism from the empty set to $M_{1}:=\alpha^{-1}(t)$ and $W_{2}$ a bordism from $M_{1}$ to $\partial W$. Of course, $W_{2}$ has only critical points $p_{i}$ with $Ind_{\alpha}(p_{i})\geq 3$. Consider now the function $-\alpha$: this is a Morse function on $W_{2}$ seen as a bordism from $\partial W$ to $M_{1}$ with same critical points $p_{i}$ but with indices now given by $Ind_{-\alpha}(p_{i})=dim(W)-Ind_{\alpha}(p_{i})$. These critical points $p_{i}$ are then associated to $G$-equivariant $(Ind_{-\alpha}(p_{i})-1)$-surgeries, hence of codimension $Index_{\alpha}(p_{i})+1$ which is $\geq 3$.
@@
    has a right inverse, i.e. there exists a $G$-equivariant map $h\colon Y^{(2)}\rightarrow X^{(2)}$ such that $f^{(2)}\circ h$ is homotopic to $i$. To see this, observe that the existence of such a map $h$ is guaranteed, up to $G$-homotopy, by Proposition 2.4. In fact, since $f^{(2)}$ is $2$-connected and $Y^{(2)}$ has dimension $\leq 2$, it suffices to apply 2.4 with $A=Y^{(2)}$, $B=X^{(2)}$, $C=Y$ and $f=f^{(2)}$ to the map $i\in[A,C]^{G}$. The surjectivity of $f_{*}$ then implies the existence of $h$.
@@
[…]
@@
Injectivity. In order to prove the injectivity of $f_{*}\colon R_{n}^{spin}(X)^{G}\to R_{n}^{spin}(Y)^{G}$, let us consider a class $[W,\varphi:W\rightarrow X,g]\in R_{n}^{spin}(X)^{G}$ such that its image $f_{*}[W,\varphi:W\rightarrow X,g]$ is equal to the trivial element in $R_{n}^{spin}(Y)^{G}$. This means that there exists
  • a $n+1$-dimensional $G$-manifold with corners $B$, whose codimension 1 faces are $W$ itself together with a bordism $V$ from $\partial W$ to the empty set, which intersect in the only codimension 2 face $\partial W=W\cap V$;
@@
By construction, the bordism $B_{2}$ is the trace of surgeries of codimension $\geq 3$. Therefore, we can apply the equivariant version [Hanke, Theorem 2] of the Gromov-Lawson Theorem to extend the metric $g_{\bar{V}}$ to a $G$-invariant metric of positive scalar curvature $g_{2}$ on $B_{2}$. Let us denote by $g_{1}$ the $G$invariant metric of positive scalar curvature obtained by restricting $g_{2}$ to $W_{1}$.
The last fact to prove is that $\Psi_{|B_{1}}\colon B_{1}\to Y$ factors through $f\colon X\to Y$. Indeed $B_{1}$ is obtained form $W$ by attaching, up to homotopy, cells of dimension up to 2. Define a map $h\colon B_{1}\times\{0\}\cup W\times[0,1]\to Y$ as the restriction of $\Psi|_{B_{1}}\circ pr$ to this subspace, where $pr\colon B_{1}\times[0,1]\to B_{1}$ is the obvious projection. Note that by assumption the restriction of this map to $W\times\{1\}$ equals $f\circ\varphi$. These are exactly the conditions of the relative precursor [ttd, (2.5) Proposition] of Proposition 2.4 which now implies in particular (because $f$ is $2$-connected) that there exists an extension $K\colon B_{1}\to X$ of $\varphi$.
@@ 5 Universal Spaces
In algebraic topology, a convenient way to deal with fundamental group information is by using the classifying space $B\Gamma$ of a (discrete) group $\Gamma$. Let us paraphrase its relevant properties:
  • it encodes precisely the fundamental group information of $\Gamma$, but not more (all other homotopy groups vanish);
  • it has the following universal property: for every connected CW-complex $X$ with basepoint $x_{0}$ the fundamental group functor gives a bijection
    $[(X,x_{0}),(B\Gamma,y_{0})]\to\operatorname{Hom}(\pi_{1}(X,x_{0}),\Gamma)$
@@
As a first step, we define the fundamental groupoid functor of a $G$-CW-complex as a natural generalization of the fundamental group of a CW-complex.
@@
whose objects are all subgroups in $\mathcal{F}$ and morphisms are $G$-maps $G/H\to G/K$. Note that any such map is of the form $xH\mapsto xgK$ for a well defined coset $[g]\in G/K$, where we also have that $g^{-1}Hg\subset K$ —precisely the condition for this map to be well defined.
@@
  • to $H\in\mathcal{I}(X)$ the fundamental groupoid of $X^{H}$ restricted to the 00-skeleton of $X^{H}$, which we denote by $\Pi_{1}(X^{H})_{|X^{H}_{(0)}}$,
  • to a morphism $G/H\mapsto G/K;xH\mapsto xgK$ between two objects $H$ and $K$ in $\operatorname{Orb}(G,\mathcal{I}(X))$ the induced homomorphism of groupoids between $\Pi_{1}(X^{K})_{|X^{K}_{(0)}}$ and $\Pi_{1}(X^{H})_{|X^{H}_{(0)}}$, induced by the map $X^{K}\to X^{H};x\mapsto gx$.
@@
assigning to $H\notin\mathcal{I}(X)$ the empty groupoid. Note that the empty groupoid has a unique map to any other groupoid, but is not the target of any map from a non-empty groupoid. This is no problem because families are closed under taking subgroups and therefore if $K\in\mathcal{I}(X)$ and $H\notin\mathcal{I}(X)$ there never is a morphism from $K$ to $H$ in $\mathcal{ALL}$.
@@
More conceptually, to construct the fundamental groupoid functor, we can observe that we have the canonical homomorphism
@@ 5.1 Fundamental groupoid functor realization
Step 1. The 0-skeleton of any $G$-CW-complex $X$ with fundamental groupoid functor $\Pi$ is directly determined by $\Pi$ itself, more precisely by the units of $\Pi(H)$ for the subgroups $H$ of $G$. Considering an orbit $G/H$ as a discrete topological space with $G$-action, the “identity” functor defines a covariant functor $E\colon\operatorname{Orb}\to\operatorname{G-TOP}$ sending $G/H$ to the discrete topological $G$-space $G/H$. Then we define (and are required to do so)
@@
Concretely, this is the (discrete) $G$-space obtained as quotient space
$\bigsqcup_{H\in\mathcal{ALL}}G/H\times\Pi(H)_{0}/\sim$ (5.1)
@@
Step 2. Let us show that this produces the desired 00-skeleton of the space $X$ to construct. For this, we have to compute the fixed point sets say for the subgroup $H$ of $G$. It is straightforward to see that $gK\in G/K$ is fixed by $H$ if and only if $g^{-1}Hg\subset K$. At the same time, from such an element $gK$, we then get a well defined $G$-map
@@
which sends the coset $1H\in G/H$ to $gK\in G/K$. Thus, all the points $(gK,x)\in G/K\times\Pi(K)_{0}$ in $X^{(0)}$ which are fixed by $H$ are identified with a point in the single copy $\{1H\}\times\Pi(H)_{0}$ and therefore the $H$-fixed set of $X^{(0)}$ is a quotient of $\Pi(H)_{0}$. We are done once we have shown that no further identifications occur. This follows from the functoriality of $\Pi$: whenever two $H$-fixed points are identified with each other, they are also identified with a single point in $\{1H\}\times\Pi(H)_{0}$.
@@
Note, as a remark, that we can not glue together these spaces in the same way as we glued together the 0-skeleta in (5.1) to produce the desired $G$-space as the gluing process could destroy the fundamental groups of the smaller fixed point sets. Instead, we have to carry out a homotopy version of the co-end construction leading to the correct 00-skeleton to obtain the following $G$-space $W$:
$\left(\bigsqcup_{H\in\operatorname{Orb}(G)}G/H\times|\Pi(H)|\sqcup\bigsqcup_{\alpha\colon G/H\to G/K\atop\in\operatorname{Hom}_{\operatorname{Orb}(G)}(H,K)}G/H\times[0,1]\times|\Pi(K)|\right)\Big{/}\sim$ (5.2)
where the equivalence relation is now generated by a multiple mapping cylinder construction, gluing the ends of the spaces associated to morphisms appropriately to the spaces associated to the objects. Concretely, this is done as follows: let us use the notation $(gH,t,x)_{\alpha}$ to denote an element of the summand $G/H\times[0,1]\times|\Pi(K)|$ associated to $\alpha\colon G/H\to G/K\in\operatorname{Hom}_{\operatorname{Orb}(G)}$. Then we declare for all morphisms $\alpha\in\operatorname{Orb}(G)$, say $\alpha\colon G/H\to G/K$
@@
$W_{0}:=\left(\bigsqcup_{H\in\operatorname{Orb}(G)}G/H\times\Pi(H)_{0}\sqcup\bigsqcup_{\alpha\colon G/H\to G/K\atop\in\operatorname{Hom}_{\operatorname{Orb}(G)}(H,K)}G/H\times[0,1]\times\Pi(K)_{0}\right)\Big{/}\sim$
@@
[…]
@@
[…]
where we glue the union of the $G/H\times|\Pi(H)|$ along their 00-skeleta with identifications. Observe that we obtain the same space by simply attaching the spaces $|\Pi(H)|$ to the already constructed 00-skeleton $X^{(0)}$ as the following push-out
[…]
In particular it is clear that the 00-skeleton of $X$ is the 00-skeleton of $Z$, i.e. precisely $X^{(0)}$.
Observe that all $1$-cells of $X$ are already contained in the subspace $Z$. This space $Z$ is obtained from a disjoint union by identifying along 00-cells. The fundamental groupoids of $X$ and of the subspaces $X^{H}$, for $H$ subgroup of $G$, are then generated by the arrows given by $1$-cells, which are all contained in the constituent subspaces of $Z$. By the van Kampen theorem, the relations in the fundamental groupoid are then generated precisely by the $2$-cells. There are two types of $2$-cells: first the $2$-cells contained in the constituent subspaces of $Z$, giving rise to the fundamental groupoids $\Pi(H)$ of $|\Pi(H)|$ and second $2$-cells $\{gH\}\times[0,1]\times c_{\alpha}$ for $\alpha\colon G/H\to G/K$ and $1$-cells $c$ in $|\Pi(K)|$.
@@
[…]
@@
To summarize, we have constructed the $G$-CW-complex $X$ such that $\Pi_{1}(X;G)=\Pi$ exactly as required.

### v2

@@ 4 Invariance of R-groups under 2-equivalence
Let $f\colon X\rightarrow Y$ be a continuous, 2-connected $G$-map between $G$-CW-complexes (in particular, it induces a bijection between components of fixed point sets and an isomorphism of the fundamental groups of all components of fixed point sets of $X$ and the corresponding ones of $Y$). Then for $\geq 6$ the functorially induced map
@@
Let us consider a class $W,\varphi\colon W\rightarrow Y,g]\in R_{n}^{spin}(Y)^{G}$. We want to find a bordant cycle whose reference map factors through $f$.
Consider W as a bordism between its boundary $\partial W$ and the empty set and choose a $G$-invariant Morse function $\alpha\colon W\to\mathbb{R}$ on it with critical points rearranged as described in [Milnor, Theorem 4.8], namely for any critical points $p_{i}$ and $p_{j}$ such that $f(p_{i})<f(p_{j})$, we have that $Ind_{f}(p_{i})<Ind_{f}(p_{j})$, where $Ind_{f}(p)$ denotes the Morse index at $p$ of the function $f$. Notice that we are going to use the enhanced version of this result to the equivariant setting, see for instance [Mayer] or [Wasserman].
Then there exists a suitable $t\in\mathbb{R}$ such that the subset $W_{1}:=\alpha^{-1}([0,t])\subset W$ consists only of $G$-handles of dimensions 0,1 and 2. We immediately obtain a decomposition of $W$ as $W_{1}\cup W_{2}$ such that $W_{1}$ is a bordism from the empty set to $M_{1}:=\alpha^{-1}(t)$ and $W_{2}$ a bordism from $M_{1}$ to $\partial W$. Of course, $W_{2}$ has only critical points $p_{i}$ with $Ind_{\alpha}(p_{i})\geq 3$. Consider now the function $\alpha$: this is a Morse function on $W_{2}$ seen as a bordism from $\partial W$ to $M_{1}$ with same critical points $p_{i}$ but with indices now given by $Ind_{-\alpha}(p_{i})=dim(W)-Ind_{\alpha}(p_{i})$. These critical points $p_{i}$ are then associated to $G$-equivariant $(Ind_{-\alpha}(p_{i})-1)$-surgeries, hence of codimension $Index_{\alpha}(p_{i})+1$ which is $\geq 3$.
@@
    has a right inverse, i.e. there exists a $G$-equivariant map $h\colon Y^{(2)}\rightarrow X^{(2)}$ such that $f^{(2)}\circ h$ is homotopic to $i$. To see this, observe that the existence of such a map $h$ is guaranteed, up to $G$-homotopy, by Proposition 2.7. In fact, since $f^{(2)}$ is $2$-connected and $Y^{(2)}$ has dimension $\leq 2$, it suffices to apply 2.7 with $A=Y^{(2)}$, $B=X^{(2)}$, $C=Y$ and $f=f^{(2)}$ to the map $i\in[A,C]^{G}$. The surjectivity of $f_{*}$ then implies the existence of $h$.
@@
${W_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi_{1}}$$\scriptstyle{\varphi_{1}}$${Y}$${Y^{(2)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$$\scriptstyle{i}$${X^{(2)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{(2)}}$$\scriptstyle{j}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$
@@
Injectivity. In order to prove the injectivity of $f_{*}\colon R_{n}^{spin}(X)^{G}\to R_{n}^{spin}(Y)^{G}$, let us consider a class $W,\varphi:W\rightarrow X,g]\in R_{n}^{spin}(X)^{G}$ such that its image $f_{*}[W,\varphi:W\rightarrow X,g]$ is equal to the trivial element in $R_{n}^{spin}(Y)^{G}$. This means that there exists
  • an $n+1$-dimensional $G$-manifold with corners $B$, whose codimension 1 faces are $W$ itself together with a bordism $V$ from $\partial W$ to the empty set, which intersect in the only codimension 2 face $\partial W=W\cap V$;
@@
By construction, the bordism $B_{2}$ is the trace of surgeries of codimension $\geq 3$. Therefore, we can apply the equivariant version [Hanke, Theorem 2] of the Gromov-Lawson Theorem to extend the metric $g_{\bar{V}}$ to a $G$-invariant metric of positive scalar curvature $g_{2}$ on $B_{2}$. Let us denote by $g_{1}$ the $G$ -invariant metric of positive scalar curvature obtained by restricting $g_{2}$ to $W_{1}$.
The last fact to prove is that $\Psi_{|B_{1}}\colon B_{1}\to Y$ factors through $f\colon X\to Y$. Indeed $B_{1}$ is obtained form $W$ by attaching, up to homotopy, cells of dimension up to 2. Define a map $h\colon B_{1}\times\{0\}\cup W\times[0,1]\to Y$ as the restriction of $\Psi|_{B_{1}}\circ pr$ to this subspace of $B_{1}\times[0,1]$, where $pr\colon B_{1}\times[0,1]\to B_{1}$ is the obvious projection. Note that by assumption the restriction of this map to $W\times\{1\}$ equals $f\circ\varphi$. These are exactly the conditions of the relative precursor [ttd, (2.5) Proposition] of Proposition 2.7 which now implies in particular (because $f$ is $2$-connected) that there exists an extension $K\colon B_{1}\to X$ of $\varphi$.
@@ 5 Universal Spaces
In algebraic topology, a convenient way to deal with fundamental group information is by using the classifying space $B\Gamma$ of a (discrete) group $\Gamma$. Let us paraphrase its relevant property:
$B\Gamma$ has the following universal property: for every connected CW-complex $X$ with basepoint $x_{0}$ the fundamental group functor gives a bijection
$(X,x_{0}),(B\Gamma,y_{0})]\to\operatorname{Hom}(\pi_{1}(X,x_{0}),\Gamma)$
@@
It is a known fact that $B\Gamma$ is characterized by the fact that it is connected, has fundamental group $\Gamma$, whereas all higher homotopy groups are trivial.
@@
As a first step, we define the fundamental groupoid functor of a $G$-CW-complex as a natural generalization of the fundamental group of a CW-complex. In order to do that, let us first recall the definition of fundamental groupoid $\Pi_{1}(X)$ of a topological space $X$: it is the groupoid whose objects are the points of $X$; whose arrows, from $x$ to $y$ for instance, are equivalence classes of paths starting at $x$ and ending at $y$, where the equivalence relation is given by homotopy of paths with fixed starting and ending points; the composition is given by the concatenation of paths. Observe that the set of arrows starting and ending at a same point $x\in X$ is just the fundamental group of $X$ at $x$.
@@
whose objects are all subgroups in $\mathcal{F}$ and morphisms from $H\in\mathcal{F}$ to $K\in\mathcal{F}$ are $G$-maps $G/H\to G/K$ for $H,K\in\mathcal{F}$. Note [^2] To be explicit: as $G/H$ consists of a single $G$-orbit, the map is determined by the image of the coset $H$. Say this image is the coset $gK$. Then by $G$-equivariance for each $x\in G$ we must have that $xH$ is mapped to $xgK$. To be well defined, for every $x\in G$ and $h\in H$ we require that $xhgK$ and $xgK$ are in the same $K$-coset, i.e. $g^{-1}x^{-2}xhg\in K$ for each $h\in H$, which is the condition $g^{-1}Hg\subset K$. that any such map is of the form $xH\mapsto xgK$ for a well defined coset $g]\in G/K$, where we also have that $g^{-1}Hg\subset K$ —precisely the condition for this map to be well defined.
@@
  • to a group $H\in\mathcal{I}(X)$ (the isotropy family) the fundamental groupoid of $X^{H}$ restricted to the 0-skeleton of $X^{H}$ (meaning that we take the full subgroupoid), which we denote by $\Pi_{1}(X^{H})_{|X^{H}_{(0)}}$,
  • to a morphism from $H$ to $K$ in the category $\operatorname{Orb}(G,\mathcal{I}(X))$ given as $G$-map $G/H\mapsto G/K;xH\mapsto xgK$ the morphism of groupoids between $\Pi_{1}(X^{K})_{|X^{K}_{(0)}}$ and $\Pi_{1}(X^{H})_{|X^{H}_{(0)}}$ induced by the map $X^{K}\to X^{H}$ defined as $x\mapsto gx$.
@@
assigning to $H\notin\mathcal{I}(X)$ the empty groupoid. Note that the empty groupoid has a unique map to any other groupoid, but is not the target of any map from a non-empty groupoid. The former is good because it determines the value of $\Pi_{1}(X;G)$ on morphisms in $\operatorname{Orb}(G)$ with domain $H$. The latter is no problem because families are closed under conjugation and under taking subgroups and therefore if $K\in\mathcal{I}(X)$ and $H\notin\mathcal{I}(X)$ there never is a morphism from $K$ to $H$ in $\mathcal{ALL}$333recall that a morphism from $H$ to $K$ exists if and only if $\exists g\in G$ with $g^{-1}Hg\subset K$.
@@
More conceptually, to construct the fundamental groupoid functor, we can observe that we have the canonical morphism
@@ 5.1 Fundamental groupoid functor realization
Step 1. The 0-skeleton of any $G$-CW-complex $X$ with fundamental groupoid functor $\Pi$ is directly determined by $\Pi$ itself, more precisely by the units of $\Pi(H)$ for the subgroups $H$ of $G$. Considering an orbit $G/H$ as a discrete topological space with $G$-action, the “identity” functor defines a covariant functor $E\colon\operatorname{Orb}(G)\to\operatorname{G-TOP}$ sending $G/H$ to the discrete topological $G$-space $G/H$. Then we define (and are required to do so)
@@
Concretely, by definition of $\otimes_{\operatorname{Orb}}$ in [DavisLueck], this is the (discrete) $G$-space obtained as quotient space
$\bigsqcup_{H\in Ob(\operatorname{Orb}(G))}G/H\times\Pi(H)_{0}/\sim$ (5.1)
@@
Step 2. Let us show that this produces the desired 0-skeleton of the space $X$ to construct. For this, we have to compute the fixed point sets say for the subgroup $H$ of $G$. It is straightforward to see that $gK\in G/K$ is fixed by $H$ if and only if $g^{-1}Hg\subset K$. At the same time, from such an element $gK$, we then get a well defined $G$-map
@@
which sends the coset $1H\in G/H$ to $gK\in G/K$, where $1$ denotes the unit in $G$. Thus, all the points $(gK,x)\in G/K\times\Pi(K)_{0}$ in $X^{(0)}$ which are fixed by $H$ are identified with a point in the single copy $\{1H\}\times\Pi(H)_{0}$ and therefore the $H$-fixed set of $X^{(0)}$ is a quotient of $\Pi(H)_{0}$. We are done once we have shown that no further identifications occur. This follows from the functoriality of $\Pi$: whenever two $H$-fixed points are identified with each other, they are also identified with a single point in $\{1H\}\times\Pi(H)_{0}$.
@@
Note, as a remark, that we can not glue together these spaces in the same way as we glued together the 0-skeleta in (5.1) to produce the desired $G$-space as the gluing process could destroy the fundamental groups of the smaller fixed point sets. Instead, we have to carry out a homotopy version of the co-end construction which has to the correct 0-skeleton to obtain the following $G$-space $W$:
$\left(\bigsqcup_{H\in Ob(\operatorname{Orb}(G))}G/H\times|\Pi(H)|\sqcup\bigsqcup_{\alpha\colon G/H\to G/K\atop\in\operatorname{Hom}_{\operatorname{Orb}(G)}(H,K)}G/H\times[0,1]\times|\Pi(K)|\right)\Big/\sim$ (5.2)
where the equivalence relation is now generated by a multiple mapping cylinder construction, gluing the ends of the spaces associated to morphisms appropriately to the spaces associated to the objects. Concretely, this is done as follows: let us use the notation $(gH,t,x)_{\alpha}$ to denote an element of the summand $G/H\times[0,1]\times|\Pi(K)|$ associated to $\alpha\colon G/H\to G/K\in\operatorname{Hom}_{\operatorname{Orb}(G)}$. Then we declare for all morphisms $\alpha\in Mor(\operatorname{Orb}(G))$, say $\alpha\colon G/H\to G/K$
@@
$W_{0}:=\left(\bigsqcup_{H\in\atop Ob(\operatorname{Orb}(G))}G/H\times\Pi(H)_{0}\sqcup\bigsqcup_{\alpha\colon G/H\to G/K\atop\in\operatorname{Hom}_{\operatorname{Orb}(G)}(H,K)}G/H\times[0,1]\times\Pi(K)_{0}\right)\Big/\sim$
@@
${W_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${W\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X^{(0)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X}$
@@
${W_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${W_{0}\cup\left(\displaystyle\bigsqcup_{H\in\operatorname{Orb}(G)}G/H\times|\Pi(H)|\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X^{(0)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Z}$
where we glue the union of the $G/H\times|\Pi(H)|$ along their 0-skeleta with identifications. Observe that we obtain the same space by simply attaching the spaces $|\Pi(H)|$ to the already constructed 0-skeleton $X^{(0)}$ as the following push-out
${\displaystyle\bigsqcup_{H\in\mathcal{ALL}}G/H\times|\Pi(H)|^{(0)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${\displaystyle\bigsqcup_{H\in\mathcal{ALL}}G/H\times|\Pi(H)|\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X^{(0)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Z}$
In particular it is clear that the 0-skeleton of $X$ is the 0-skeleton of $Z$, i.e. precisely $X^{(0)}$.
Observe that all $1$-cells of $X$ are already contained in the subspace $Z$. This space $Z$ is obtained from a disjoint union by identifying along 0-cells. The fundamental groupoids of $X$ and of the subspaces $X^{H}$, for $H$ subgroup of $G$, are then generated by the arrows given by $1$-cells, which are all contained in the constituent subspaces of $Z$. By the van Kampen theorem, the relations in the fundamental groupoid are then generated precisely by the $2$-cells. There are two types of $2$-cells: first the $2$-cells contained in the constituent subspaces of $Z$, giving rise to the fundamental groupoids $\Pi(H)$ of $|\Pi(H)|$ and second $2$-cells $\{gH\}\times[0,1]\times c_{\alpha}$ for $\alpha\colon G/H\to G/K$ and $1$-cells $c_{\alpha}$ in $|\Pi(K)|$.
@@
${(W_{0})^{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${(W_{0})^{H}\cup\left(\displaystyle\bigsqcup_{K\in\operatorname{Orb}(G);\atop\,\{gK\mid g^{-1}Hg\subset K\}}gK\times|\Pi(K)|\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(X^{(0)})^{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Z^{H}}$
@@
To summarize, we have constructed the $G$-CW-complex $X$ such that $\Pi_{1}(X;G)=\Pi$ exactly as required. ∎